OpenStudy (anonymous):

If Log"2n"(1944) = Log"n" (486√2), determine the value of n^6. i got 3.2134 can anyone else answer this? cause i no im wrong

OpenStudy (anonymous):

we can do this

OpenStudy (anonymous):

it is a pain, but we can do it.

OpenStudy (anonymous):

ok cool

OpenStudy (anonymous):

the "2n" and "n" are subscripts of log

OpenStudy (anonymous):

ok first of all we write these in equivalent exponential form. i will put A = 1944 and B = \[486\sqrt{2}\] and we will put the numbers back when we need them . now now

OpenStudy (anonymous):

so this says that \[(2n)^{log_{2n}(A)}=(2n)^{log_n(B)}\]

OpenStudy (anonymous):

let me know if you get stuck an any step. i just expontiated both sides, that is i took 2n to the power of both sides

OpenStudy (anonymous):

the left hand side is just A. the right hand side is \[2^{\log_n(B)}\times n^{\log_n(B)}\] so we get \[A=2^{\log_n(B)}\times n^{\log_n(B)}\]

OpenStudy (anonymous):

now \[n^{\log_n(B)}=B\] so the above equation is \[A=2^{log_n(B)}\times B\]giving \[\frac{A}{B}=2^{\log_n(B)}\]

OpenStudy (anonymous):

how we doing so far?

OpenStudy (anonymous):

im trying to write it down but still a lil lost

OpenStudy (anonymous):

well i will keep writing and then you can ask about any step. i am writing all the steps and will try to give reasons for each one

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

starting now with \[\frac{A}{B}=2^{\ln_n(B)}\] writing in equivalent log form (now the base will be 2) we get \[\log_2(\frac{A}{B})=\log_n(B)\]

OpenStudy (anonymous):

now writing in equivalent exponential form we solve for B to get \[B=n^{\log_2(\frac{A}{B})}\]

OpenStudy (anonymous):

Is it n^6 = 2.2315 * 10^1?

OpenStudy (anonymous):

actually i am not sure but lets keep going. before we go any further, lets put in the numbers

OpenStudy (anonymous):

\[\frac{A}{B}=\frac{1994}{486\sqrt{2}}=4\sqrt{2}=\sqrt{32}\]

OpenStudy (anonymous):

so we actually know what \[\log_2(\frac{A}{B}) \] is. it is \[\log_2(\sqrt{32})=\frac{5}{2}\] since \[2^{\frac{5}{2}}=\sqrt{32}\]

OpenStudy (anonymous):

hmm im lost still don't no how but im lost

OpenStudy (anonymous):

so \[n^{\frac{5}{2}}=486\sqrt{2}\] \[n=(486\sqrt{2})^{\frac{2}{5}}\] \[n^6=(486\sqrt{2})^{\frac{12}{5}}\]

OpenStudy (anonymous):

k i no where i got lost at